A rectangular park is 100yards wide and 150yards long.
Give the length and width of another rectangular park that has the same perimeter but a larger area.

Answer by richwmiller(17219)

A square would have greatest area.
100*150=15000
2*100*2*150=500
500/4=125 each side
125^2=15625